Accelerator Update for Jun 12, 2006 - June 13, 2006
Monday June 12
The day shift began with the Antiproton Source (Pbar) stacking, with the Booster, Main Injector (MI), Recycler, and Tevatron (TeV) conducting studies, with experts investigating the NuMI Horn power supply ground fault, and with MiniBooNE taking beam.
Operations reported at 10:59 AM that the NuMI Horn hadn’t tripped off for a while. A NuMI expert asked for some low intensity beam.
Operations sent beam to NuMI at 11:50 AM.
Operations put the TeV into a dry squeeze at 12:44 PM.
Recycler experts began a study period around 1 PM.
The Recycler stash ended around 6 PM.
Operations put the TeV into shot setup around 7:19 PM.
Tuesday June 13
The midnight shift began with Operations monitoring store 4759 and a small stash, with Pbar stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, and with MiniBooNE and NuMI taking beam.
At 12:47 AM, the Klystron Debuncher cavity water skid tripped off on a water flow fault for the first of twenty times in a little over five hours. It reset each time.
Recycler began preparing to stash at 4:45 AM.
The stash ended and Pbar resumed stacking at 6:29 AM.
The Plans for Tuesday
The plans for today are to leave the store in until later this afternoon, and to try and resolve the sequencer and SDA problems. Booster experts will be allowed to conduct some destructive orbit and extraction tuning. Pbar experts will be allowed to conduct target spot size measurements.
Comments and Suggestions
|last modified 06/13/2006 email Fermilab|